# DIVERSE: Deciphering Internet Views on the U.S. Military Through Video Comment Stance Analysis, A Novel Benchmark Dataset for Stance Classification

### v1

@@ Introduction
Finally, in the case of labeling the stance of the comment toward its respective video, we also used the stance of the comment toward the U.S. Army as a weak label for this stance. Generally speaking, if a comment was against the U.S. Army, it would also be against a video posted by the U.S. Army, and the use of complementary stances has been shown to improve performance in the stance detection task (Sobhani, Inkpen, and Zhu 2017; Zhang et al. 2023b). Thus, we used the final stance toward the U.S. Army produced by Data Programming on the aforementioned weak labels as a weak label for the stance toward the video. It is important to note that this relationship does not work the other way; we found several examples where commentators would express a supportive stance toward the U.S. Army but would be very against a particular video.
Determining Stance
@@
Having created the weak labels for the comments, we then use the Data Programming methodology in the form of the snorkel Python package [^7] https://snorkel.readthedocs.io/en/master/index.html, which uses a unified generative model that takes into account partial coverage of the weak labels as well as correlations between the weak labels to produce the final labels. We refer the reader to (Ratner et al. 2016) and (Ratner et al. 2019) for the specifics on the function, how it is trained, and how it produces the final labels. For the final labels, we allow the stance to take the form of “supports”, “against”, and “neutral” for both the stance toward the video and the stance toward the U.S. Army.
FAIR Aspects of Dataset
Findability & Accessibility
We provide public access to our dataset and source code at the indexed-dataset sharing service Zenodo, https://doi.org/10.5281/zenodo.10493803 to facilitate replication and use of this dataset. A meta-data file is also provided to briefly describe each of the fields. In particular, we also provide the weak labels as part of the dataset as well.
Interoperability & Reusability
The dataset is provided in a generic .csv format which supports reusability across several data analysis applications. We also provide the code used to generate the dataset for further reusability. This code includes the code used to produce the weak labels as well as the code to create the final labels from the weak labels.
Ethical Considerations
In the collection of this dataset, we only collected data provided by the YouTube API and did not collect any private profiles or videos. We anonymized the author IDs, both in the author field of the data as well as in the comments section if a comment were a reply to another comment, and we did not perform any author-level analysis.
Dataset Properties
In this section, we outline some properties of the data as well as the stance labels to understand the dataset.
Data Properties
@@
The dataset consists of 858 videos published from 1 June 2010 through 5 October 2023 (final date of collection), all from the U.S. Army’s official YouTube channel. Figure 2 depicts the distribution of the videos on the channel over time
@@
Generally speaking, the channel has published an increasing number of videos over time, with some occasional periods of fewer videos. Table 1 summarizes the statistics for the comments to each video. Most videos within the data set only have a single comment, with the number of comments per video being highly skewed and some videos having a large number of comments (the most commented-on video, ‘Soldier wakes up and chooses “HOOAH!”’, has 9.4% of all of the comments by itself). When it comes to the timing of the comments themselves, we find the overwhelming have occurred recently in the channel’s timeline (since 2022). Figure 3, depicts the distribution of comments, across all videos, by month and year.
@@
Table 1: Per Video Comment Statistics
From the plot of when comments are published, there are significant spikes in comments in August, September, and October of 2022 and again in May and June of 2023. In the case of the first spike, the most commented-on video was ‘Soldier wakes up and chooses “HOOAH!”’, while the second spike in the Spring of 2023 has ‘Words of wisdom from a #WWII veteran’ the most commented-on video. For the first video, it seems the commentary is divisively split about what the Army does or should do in the world, while in the second case the comments seem to both positively reminisce about the Army in World War II and lament about its current state. Turning to the comments themselves, we find that most comments are relatively short, like many social media posts, but a few comments are much longer at over 1,000 words for the comment. Table 2 has the word count statistics for the comments.
Word-level Statistics of Comments Value
Mean 16.77
Standard Deviation 32.67
Minimum 1.00
25th Percentile 4.00
50th Percentile (Median) 9.00
75th Percentile 18.00
Maximum 1851.00
Table 2: Word-level Statistics of the Comments
Label Properties
Turning now to stance labels for the dataset, we find that the labels skew slightly toward the “against” characterization for both the U.S. Army and the videos being posted on the channel. The following table, 3 summarizes the label counts for each of the stances.
@@
Table 3: Labels Counts (and percentages of the total number) for Each of the Stance Targets
Interestingly, from the label counts we also observe both a relatively even distribution of labels across the stances but also the “neutral” stance being the minority label. In most other benchmark stance datasets, there is a strong label skew, particularly for datasets like covid-lies (Hossain et al. 2020) and phemerumors (Kochkina, Liakata, and Zubiaga 2018), the “neutral” stance tends to be the majority stance label. Since this is a first-of-its-kind dataset for both the targets as well as the data domain (i.e., YouTube) it’s not clear if the targets tend to inspire stronger stance-taking behavior in comments or if the domain of YouTube causes stronger stance-taking behavior. Regardless, upon manual inspection of some of the labels, we do find that these stance labels are accurate and that there is a lot of strong stance-taking within the comments.
Turning to the distribution of stance labels over time, we find that each of the three classes of stance generally trend together over time. Figure 4 displays the distribution of the different classes of stance toward the U.S. Army for all of the comments (the stances toward the videos follow a very similar trend). From this figure, we can observe that the two prominent periods of comments do differ somewhat in their compositions of stance labels. In the first peak of comments, centering in September of 2022, we see more “against” stance labels than “supports” or “neutral”. Whereas in the second spike of comments around May-June of 2023, we see more “supports” comments than “against” or “neutral”. Thus, overall, there is a moderate balance among the stances expressed in the comments toward the videos, but this balance does shift over time and between videos.
@@
To better understand these differences in stances, we also looked at the videos with the most positive, neutral and negative receptions. For videos with more than 100 comments, We found the video, “Honoring Our Fallen Heroes - Taps” had the highest percentage of supportive comments at 76.5% of the comments supporting the video. On the opposite end was the video “What does the Army of the future look like?” which features a short interview with former Sergeant Major Grinston on the future of the U.S. Army, which had 66.9% of comments that were against the video. And, the most neutrally received video was “Army #PopQuiz - What training is this?” which received 66.5% of its comments as being neutral toward the video, many of them answering the question posed in the video.
@@
Leave-One-Out Stance Detection
To further validate the stance labels produced by the proposed Data Programming method, we performed a leave-one-out, out-of-domain experiment referenced from (Ng and Carley 2022). We trained a model on the stance labels and then evaluated that model on the other benchmark datasets. In our test, we trained a cased BERT model [^8] https://huggingface.co/bert-base-cased (Devlin et al. 2018) on only the stance labels from the DISMISS dataset and then evaluated the model on all of the benchmark datasets. The model was trained for 3 epochs using a learning rate of $41e-$ 4 with linear decay, a batch size of 32, and was trained to classify text into three different labels. The text that was input to the model was given as:
@@
In this test, if the proposed data labels are capturing the stances correctly, then we should see similar performance on the benchmark datasets when predicting stance, as was seen in (Ng and Carley 2022). Table 4 has the out-of-domain test results. The results are evaluated using the Macro F1 score as a suitable evaluation method due to the almost equal distribution of stance labels.
On average, the model trained on stances only from this dataset achieves a macro F1 score of 0.3043. This is similar to the average accuracy of 0.303 in a past work that performed leave-one-out stance detection using multiple BERT-based classifiers (Ng and Carley 2022). This therefore suggests that the stance labels proposed in this paper are valid stance labels.
@@
Table 4: Macro F1 Scores for Model Applied to Different Datasets
Discussion
Your text is well-written, but I’ve made some minor corrections for clarity and consistency. Here’s the revised version:
Public opinion towards a country’s army is crucial, as it signals the level of trust in the army and the willingness of citizens to join. A positive opinion facilitates army recruitment. Social media engagement plays a pivotal role in shaping this opinion. Therefore, this paper introduces a dataset that supports such understanding and associated analyses, comprising comments on the U.S. Army’s official YouTube videos annotated for their stance towards the Army and the video content itself.
To construct this dataset, we employed the Data Programming paradigm, which leverages weak labels for automated labeling. In this paradigm, humans transition from direct annotation to machine-assisted annotation, allowing scalability to larger datasets. For this specific dataset, we utilized subject matter expertise, domain knowledge, and recent advancements in using LLMs for stance determination to create weak labels. Specifically, we used weak labels such as the presence of hate speech, sarcasm, and sentiment as indicators for certain stances. This transformation converted the challenging, ambiguously defined problem of determining and labeling stance into a more well-defined problem for specific aspects or stances. These could then be aggregated through Data Programming to produce comprehensive stance labeling. For instance, the presence of hate speech indicates a comment against the video and, by extension, against the user who uploaded it (e.g., the Army).
We further enhanced these weak labels with contributions from LLMs. Currently, the effectiveness of LLMs in tasks like stance labeling remains uncertain. While recent research shows promise for using LLMs, especially when combined with prompt engineering, for stance detection and labeling, the generalizability of these methods and their compatibility with all LLMs is unclear. Through the Data Programming label, we treat these LLM-derived labels as noisy yet valuable, enhancing the creation of a stance-labeled dataset. The combination of dataset-specific weak signals and the linguistic understanding power of LLMs produces high-quality labels for stance in a more scalable manner compared to purely manual or machine-generated labels.
@@ Discussion
The dataset proposed in this work has several potential uses. First, this dataset presents a versatile resource for research focusing on the analysis of stance in YouTube video comments. Beyond its primary application in training and evaluating new stance detection techniques and models, the dataset introduces distinctive traits not present in current benchmarks. Notably, it originates from YouTube, offering a departure from the prevalent Twitter-based datasets. Furthermore, the dataset features multiple stance targets and exhibits a substantial degree of stance-taking behavior and controversial content (i.e., a high proportion of “supports” and “against” stances). The dataset provides sequential information about comments, allowing investigations into the dynamics of stance with respect to the stance expressed in a previous comment.
@@
Given the dataset’s proximity to a politically and culturally divisive topic — the U.S. Military — it encompasses a spectrum of online behaviors, including the propagation of conspiracy theories, trolling, hate speech, and misinformation. As such, it serves as a valuable resource for studies on misinformation and disinformation.
@@
Moreover, the dataset’s temporal collection allows researchers to explore dynamic aspects of stance-taking behavior over time. This temporal dimension provides insights into the evolution of discussions and sentiments within the context of the U.S. Army’s YouTube channel. In summary, the dataset offers a rich foundation for various studies, ranging from advancing stance detection methodologies to delving into the nuanced dynamics of online interactions related to a prominent and controversial subject.
Limitations
@@
Our dataset is limited to the comments focusing on the US Army, and only those we could collect. Not all of the videos posted to the channel allowed comments. We hope to expand our dataset to include the comments towards videos posted by armies around the world, to evaluate public opinion towards the army. In addition, our dataset is limited to the stance annotation of local, relatively small LLMs due to the size of the dataset and computational constraints. In the future, we hope to expand the dataset annotation to different types of LLMs with varying prompting schemes to better understand the stance annotation capabilities of LLMs.
Conclusion
@@
Understanding the stance expressed in YouTube comments towards an entity is essential for gaining insights into public opinion regarding the authors of the videos. In this context, we introduce the DIVERSE dataset, a dataset that captures the stance towards the US Army. This dataset is formulated from comments gathered from the U.S. Army’s YouTube videos. We constructed the final stance variable utilizing weak labels, including the identification of hate speech and sarcasm, and incorporated stance inferences from Large Language Models. By employing a data programming model, we amalgamated these diverse labels to generate the final stance annotations. We anticipate that this dataset will prove invaluable to researchers investigating public sentiments and stances toward military entities.
Acknowledgments
@@
The research for this paper was supported in part by the Center for Informed Democracy and Social-cybersecurity (IDeaS) and the Center for Computational Analysis of Social and Organizational Systems (CASOS) at Carnegie Mellon University. This work was also conducted within the Cognitive Security Research Lab at the Army Cyber Institute. The views and conclusions are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Department of Defense, the U.S. Army, or the U.S. Government.
Appendix A Appendix 1: Ethics Paper Checklist
  1. 1.
    For most authors…
    1. (a)
      Would answering this research question advance science without violating social contracts, such as violating privacy norms, perpetuating unfair profiling, exacerbating the socio-economic divide, or implying disrespect to societies or cultures? Yes
    2. (b)
      Do your main claims in the abstract and introduction accurately reflect the paper’s contributions and scope? Yes
    3. (c)
      Do you clarify how the proposed methodological approach is appropriate for the claims made? Yes
    4. (d)
      Do you clarify what are possible artifacts in the data used, given population-specific distributions? We discuss the limitations of the dataset as well as give basic statistics on the dataset to clarify population-level distributions.
    5. (e)
      Did you describe the limitations of your work? Limitations are described in the limitations subsection in the discussion section.
    6. (f)
      Did you discuss any potential negative societal impacts of your work? Yes
    7. (g)
      Did you discuss any potential misuse of your work? Yes
    8. (h)
      Did you describe steps taken to prevent or mitigate potential negative outcomes of the research, such as data and model documentation, data anonymization, responsible release, access control, and the reproducibility of findings? Yes. We follow standard anonymization procedures for the social media data
    9. (i)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? Yes
  2. 2.
    Additionally, if your study involves hypotheses testing…
    1. (a)
      Did you clearly state the assumptions underlying all theoretical results? NA
    2. (b)
      Have you provided justifications for all theoretical results? NA
    3. (c)
      Did you discuss competing hypotheses or theories that might challenge or complement your theoretical results? NA
    4. (d)
      Have you considered alternative mechanisms or explanations that might account for the same outcomes observed in your study? NA
    5. (e)
      Did you address potential biases or limitations in your theoretical framework? NA
    6. (f)
      Have you related your theoretical results to the existing literature in social science? NA
    7. (g)
      Did you discuss the implications of your theoretical results for policy, practice, or further research in the social science domain? NA
  3. 3.
    Additionally, if you are including theoretical proofs…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? NA
    2. (b)
      Did you include complete proofs of all theoretical results? NA
  4. 4.
    Additionally, if you ran machine learning experiments…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? We include all of the code in a URL
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? Yes.
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? NA
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? Yes.
    5. (e)
      Do you justify how the proposed evaluation is sufficient and appropriate to the claims made? Yes.
    6. (f)
      Do you discuss what is “the cost“ of misclassification and fault (in)tolerance? Yes.
  5. 5.
    Additionally, if you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? NA
    2. (b)
      Did you mention the license of the assets? NA
    3. (c)
      Did you include any new assets in the supplemental material or as a URL? NA
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? NA. All data is publicly available data. We were also asked to analyze this data by the organization that owns the YouTube channel from which the data comes from.
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? Yes. We mention the data contains various types of conspiracy theories, hate speech ad misinformation. We also anonymize the data to remove user names to and within comments
    6. (f)
      If you are curating or releasing new datasets, did you discuss how you intend to make your datasets FAIR (see fair)? Yes.
    7. (g)
      If you are curating or releasing new datasets, did you create a Datasheet for the Dataset (see gebru2021datasheets)? Yes.
  6. 6.
    Additionally, if you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots? NA
    2. (b)
      Did you describe any potential participant risks, with mentions of Institutional Review Board (IRB) approvals? NA
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? NA
    4. (d)
      Did you discuss how data is stored, shared, and deidentified? Yes

### v2

@@ III-C6 Use of Analogous Stances
Finally, in the case of labeling the stance of the comment toward its respective video, we also used an analogous stance for a weak label. We used the stance of the comment toward the US Army. Generally speaking, if a comment was against the US Army, it would also be against a video posted by the US Army, and the use of complementary stances has been shown to improve performance in the stance detection task Sobhani et al. (2017); Zhang et al. (2023b). Thus, we used the final stance toward the US Army produced by weak supervision on the aforementioned weak labels as a weak label for the stance toward the video. It is important to note that this relationship does not work the other way; we found several examples where commentators would express a supportive stance toward the US Army but would be very against a particular video.
III-D Final Stance Determination
@@
Having created the weak labels for the comments, we then use the weak supervision methodology in the form of the snorkel Python package [^8] https://snorkel.readthedocs.io/en/master/index.html, which uses a unified generative model that takes into account partial coverage of the weak labels as well as correlations between the weak labels to produce the final labels. We refer the reader to Ratner et al. (2016) and Ratner et al. (2019) for the specifics on the function, how it is trained, and how it produces the final labels. For the final labels, we allow the stance to take the form of “supports”, “against”, and “neutral” for both the stance toward the video and the stance toward the US Army.
III-E Human evaluation
After machine annotation, we had four individuals annotate how many? what percentage of total? comments. For each comment, the authors independently annotated the stance towards the US Army and the stance towards the video. The manual annotations were then analyzed for inter-annotator agreement, and the machine annotations were compared against the manual ones.
IV Data Properties and Analysis
@@
This section describes some exploratory data analysis performed on this dataset. The dataset consists of 858 videos published from 1 June 2010 through 5 October 2023 (final date of collection) on the US Army’s official YouTube Channel.
IV-A Dataset Properties
Distribution of videos over time
@@ IV-A Dataset Properties
Generally speaking, the channel has published an increasing number of videos over time, with some occasional periods of fewer videos. Figure 3 depicts the distribution of all the videos on the channel over time. Table II summarizes the statistics for the comments to each video. Most videos within the data set only have a single comment, with the number of comments per video being highly skewed and some videos having a large number of comments (the most commented-on video, ‘Soldier wakes up and chooses “HOOAH!”’, has 9.4% of all of the comments by itself). When it comes to the timing of the comments themselves, we find the overwhelming have occurred recently in the channel’s timeline (since 2022).
Figure 2: Distribution of comments over time
From Figure 2, the plot of comments over time, there are significant spikes in comments in August, September, and October of 2022 and again in May and June of 2023. In the case of the first spike, the most commented-on video was ‘Soldier wakes up and chooses “HOOAH!”’, while the second spike in the Spring of 2023 has ‘Words of wisdom from a #WWII veteran’ the most commented-on video. For the first video, it seems the commentary is divisively split about what the Army does or should do in the world, while in the second case the comments seem to both positively reminisce about the Army in World War II and lament about its current state. Turning to the comments themselves, we find that most comments are relatively short, like many social media posts, but a few comments are much longer at over 1,000 words for the comment. Table I has the word count statistics for the comments.
Word-level Statistics of Comments Value
Mean 16.77
Standard Deviation 32.67
Minimum 1.00
25th Percentile 4.00
50th Percentile (Median) 9.00
75th Percentile 18.00
Maximum 1851.00
TABLE I: Word-level Statistics of the Comments
@@
TABLE II: Per Video Comment Statistics
Figure 3: Distribution of videos published over time
IV-B Stance Label Properties
Turning now to stance labels produced through our methodology for the dataset, we find that the labels skew slightly toward the “against” characterization for both the US Army and the videos being posted on the channel. Table III summarizes the label counts for each of the stances.
Interestingly, from the label counts we also observe both a relatively even distribution of labels across the stances but also the “neutral” stance being the minority label. In most other benchmark stance datasets, there is a strong label skew, particularly for datasets like Hossain et al. (2020) and Kochkina et al. (2018), the “neutral” stance tends to be the majority stance label. Since this is a first-of-its-kind dataset for both the targets as well as the data domain (i.e., YouTube) it’s not clear if the targets tend to inspire stronger stance-taking behavior in comments or if the domain of YouTube causes stronger stance-taking behavior. Regardless, upon manual inspection of some of the labels, we do find that these stance labels are accurate and that there is a lot of strong stance-taking within the comments.
@@ IV-B Stance Label Properties
TABLE III: Labels Counts (and percentages of the total number) for Each of the Stance Targets
Turning to the distribution of stance labels over time, we find that each of the three classes of stance generally trend together over time. Figure 4 displays the distribution of the different classes of stance toward the US Army for all of the comments. The different stances toward the videos follow a very similar trend. We further observe that the two prominent periods of comments do differ somewhat in their compositions of stance labels. In the first peak of comments, centering in September of 2022, we see more “against” stance labels than “supports” or “neutral”. Whereas in the second spike of comments around May-June of 2023, we see more “supports” comments than “against” or “neutral”. Thus, overall, there is a moderate balance among the stances expressed in the comments toward the videos, but this balance does shift over time and between videos.
@@
To better understand these differences in stances, we looked at the videos with the most positive, neutral and negative receptions. Among the videos with more than 100 comments, we found the video, “Honoring Our Fallen Heroes - Taps” had the highest percentage of supportive comments at 76.5% of the comments supporting the video. On the opposite end was the video “What does the Army of the future look like?” which features a short interview with former Sergeant Major Grinston on the future of the US Army. The video had 66.9% of comments that were against it. And, the most neutrally received video was “Army #PopQuiz - What training is this?” which received 66.5% of its comments as being neutral toward the video, many of them answering the question posed in the video.
@@
IV-C Comparison with human evaluation
The inter-annotator agreement was Iain, results here. Comparing the machine-based annotations with the human-based annotations, we get an accuracy of how many. elaborate on accuracy
IV-D Leave-One-Out Stance Detection
To further validate the stance labels produced by the proposed weak supervision method, we performed a leave-one-out, out-of-domain experiment referenced from (Ng and Carley, 2022). We trained a model on the stance labels and then evaluated that model on the other benchmark datasets. In our test, we trained a cased BERT model [^9] https://huggingface.co/bert-base-cased (Devlin et al., 2018) on only the stance labels from the DISMISS dataset and then evaluated the model on all of the benchmark datasets. The model was trained for 3 epochs using a learning rate of $41e-$ 4 with linear decay, a batch size of 32, and was trained to classify text into three different labels. The text that was input to the model was given as:
@@ IV-D Leave-One-Out Stance Detection
In this test, if the proposed data labels are capturing the stances correctly, then we should see similar performance on the benchmark datasets when predicting stance, as was seen in Ng and Carley (2022). Table IV has the out-of-domain test results. The results are evaluated using the Macro F1 score as a suitable evaluation method due to the almost equal distribution of stance labels.
On average, the model trained on stances only from this dataset achieves a macro F1 score of 0.3043. This is similar to the average accuracy of 0.303 in a past work that performed leave-one-out stance detection using multiple BERT-based classifiers Ng and Carley (2022). This therefore suggests that the stance labels proposed in this paper are valid stance labels.
@@
TABLE IV: Macro F1 Scores for Model Applied to Different Datasets
V Discussion
Public opinion towards a country’s army is crucial, as it signals the level of trust in the army and the willingness of citizens to join the forces. A positive opinion facilitates army recruitment and support, especially in times of crisis. Social media engagement plays a pivotal role in shaping this opinion. This paper introduces a dataset that supports such understanding and associated analyses, comprising comments on the US Army’s official YouTube videos annotated for their stance towards the Army and the video content itself.
To construct this dataset, we employed weak supervision, which leverages weak labels for automated labeling. This paradigm provides a transition from direct human annotation to machine-assisted annotation, allowing scalability to larger datasets. For our specific dataset, we utilized subject matter expertise, domain knowledge, and recent advancements in using LLMs for stance determination to create weak labels. Specifically, we used weak labels such as the presence of hate speech, sarcasm, and sentiment as indicators for certain stances. This transformation converted the challenging, ambiguously defined problem of determining and labeling stance into a more well-defined problem for specific aspects or stances. These could then be aggregated through weak supervision to produce comprehensive stance labeling. For instance, the presence of hate speech indicates a comment against the video and, by extension, against the user who uploaded it (e.g., the Army).
We further enhanced these weak labels with contributions from LLMs. Currently, the effectiveness of LLMs in tasks like stance labeling remains uncertain. While recent research shows promise for using LLMs, especially when combined with prompt engineering, for stance detection and labeling, the generalizability of these methods and their compatibility with all LLMs is unclear. Through weak supervision, we treat these LLM-derived labels as noisy yet valuable, enhancing the creation of a stance-labeled dataset. The combination of dataset-specific weak signals and the linguistic understanding power of LLMs produces high-quality labels for stance in a more scalable manner compared to purely manual or machine-generated labels.
@@ V Discussion
The dataset proposed in this work has several potential uses. First, this dataset presents a versatile resource for research on the analysis of stance in YouTube video comments. Beyond its primary application in training and evaluating new stance detection techniques and models, the dataset introduces distinctive traits not present in current benchmarks. Notably, it originates from YouTube, offering a departure from the prevalent X-based datasets. Furthermore, the dataset features multiple stance targets and exhibits a substantial degree of stance-taking behavior and controversial content (i.e., a high proportion of “supports” and “against” stances). The dataset provides sequential information about comments, allowing investigations into the dynamics of stance with respect to the stance expressed in a previous comment.
@@
Given the dataset’s proximity to a politically and culturally divisive topic — the military — it encompasses a spectrum of online behaviors, including the propagation of conspiracy theories, trolling, hate speech, and misinformation. As such, it serves as a valuable resource for studies on misinformation and disinformation.
@@
Finally, the dataset’s temporal collection allows researchers to explore dynamic aspects of stance-taking behavior over time. This temporal dimension provides insights into the evolution of discussions and sentiments within the context of the US Army’s YouTube channel. In summary, the dataset offers a rich foundation for various studies, ranging from advancing stance detection methodologies to delving into the nuanced dynamics of online interactions related to a prominent and controversial subject.
Dataset Availability
We collected data provided by the YouTube API and did not perform additional scraping. We only collected public videos and comments, and made no attempt to collect private information. We anonymized the author IDs and did not perform any author-level analysis. The code used for construction of the dataset and the dataset itself is provided at https://doi.org/10.5281/zenodo.10493803.
Limitations and Future Work
@@
Our dataset is limited to the comments from one channel, and the comments we could collect (not all videos posted to the channel allowed comments). We hope to expand our dataset to include the comments towards videos posted by armies around the world, to evaluate public opinion towards the army, across the world. In addition, our dataset is limited to relatively small LLMs due to computational constraints. In the future, we hope to expand the dataset annotation to different types of LLMs with varying prompting schemes to better understand the stance annotation capabilities of LLMs.
VI Conclusion
@@
Understanding the stance expressed in YouTube comments towards an entity is essential for gaining insights into public opinion regarding the authors of the videos. Here, we introduce the DIVERSE dataset, a dataset that captures the stances towards the US Army. This dataset is formulated from comments gathered from the US Army’s YouTube videos. We constructed the final stance variable utilizing weak labels, including the identification of hate speech and sarcasm, and incorporated stance inferences from Large Language Models. By employing weak supervision, we amalgamated these diverse labels to generate the final stance annotations. We anticipate that this dataset will prove invaluable to researchers investigating public sentiments and stances toward military entities.
Acknowledgements
@@
The research for this paper was supported in part by the Center for Informed Democracy and Social-cybersecurity (IDeaS) at Carnegie Mellon University. This work was also conducted within the Cognitive Security Research Lab at the Army Cyber Institute. The views and conclusions are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Department of Defense, the U.S. Army, or the U.S. Government.